# MoLA: Motion Generation and Editing with Latent Diffusion Enhanced by Adversarial Training

### v3

@@ 4. Experiments
We evaluate the performance of MoLA on two tasks: motion generation (Section 4.1) and motion editing (Section 4.2). Our results demonstrate that MoLA achieves its three key objectives: (1) fast and high-quality generation, (2) variable-length generation, and (3) multiple motion editing tasks in a training-free manner. To validate these objectives, we utilize the HumanML3D (Guo et al., 2022).333 This dataset contains 14,616 human motions from the AMASS (Mahmood et al., 2019) and HumanAct12 (Guo et al., 2020) datasets and 44,970 text descriptions. Details of the experimental setup as well as additional results on KIT-ML (Plappert et al., 2016) are provided in the Appendix.
4.1. Motion Generation
4.1.1. Performance comparison with other text-to-motion models
We evaluate generation quality and inference cost in comparison to existing text-to-motion methods that have publicly available implementations (Tevet et al., 2023; Chen et al., 2023; Zhang et al., 2023b; Pinyoanuntapong et al., 2024b; Guo et al., 2024). As shown in Figure 2, inference with our model can be achieved with equivalent time to MLD, which is much faster in generation than data-space diffusion MDM (Tevet et al., 2023) that is the only existing method that can perform training-free motion editing. It is true that most methods based on discrete representation perform very well in terms of fidelity and speed, but there is no discrete method that can deal with multiple editing tasks in a training-free manner and satisfies our requirements.
@@ 4.1.1. Performance comparison with other text-to-motion models
In addition, we evaluate our proposed MoLA in comparison to current SOTA methods (Zhang et al., 2024; Tevet et al., 2023; Wang et al., 2023; Zhong et al., 2023; Guo et al., 2024; Lou et al., 2023; Pinyoanuntapong et al., 2024b; Zhang et al., 2023b; Chen et al., 2023; Zou et al., 2024; Pinyoanuntapong et al., 2024a) using five metrics (R-Precision, Fréchet Inception Distance (FID), Multi-modal distance (MMDist), Diversity, and MultiModality (MModality)) proposed by Guo et al. (Guo et al., 2022). For evaluation, we select the model that achieves the best FID, which is a metric that evaluates the overall motion quality, on the validation set and report its performance on the test set of HumanML3D. We show the results in Table 1. The methods are organized into three groups: i) those using VQ-based latent representations (Discrete), ii) those using data-space diffusion model (Continuous (raw data)), and iii) those using VAE-based latent representations (Continuous (latent)). The discrete approaches perform well in motion generation (e.g., (Guo et al., 2024; Lou et al., 2023; Pinyoanuntapong et al., 2024b)). However, those models cannot control an arbitrary set of joints in a training-free manner (Pinyoanuntapong et al., 2024b) as we have discussed so far. MDM (Tevet et al., 2023), MotionDiffuse (Zhang et al., 2024) and Fg-T2M (Wang et al., 2023) adopt data-space diffusion, while MLD (Chen et al., 2023), MotionLCM (Dai et al., 2024) and our MoLA utilize a diffusion model in a lower-dimensional latent space. Therefore, the former are grouped in Continuous (raw data) and the latter in Continuous (latent) in the table. MoLA achieves the best performance in continuous methods, especially in the R-precision, FID, and MMDist metrics as shown in Table 1. Moreover, as shown in Figure 2, our MoLA are much faster in generation than MDM, which is the only existing method that can perform training-free motion editing.
Figure 4. Comparison of motion length distributions between the HumanML3D test set and the generated motion samples.
@@
Table 2. Analysis of motion reconstruction /generation performance on HumanML3D
Editing type Methods
R-Precision
Top-3 $\uparrow$
Overfitting-based editing OmniControl 0.688 0.192 9.533 0.065 0.007 0.053 74.4
MotionLCM 0.759 0.501 9.293 0.237 0.054 0.164 0.02
Table 3. Comparison of motion control (path-following task) on HumanML3D dataset
Next, we discuss the effects of the activation variable in our motion representation, introduced in Section 3.1. By incorporating the activation variable into the motion features in Equation (2), our stage 2 model in Section 3.3 enables the generation of variable length motions. Figure 4 shows the distribution of sample lengths generated by our model compared to the test set of HumanML3D. It is demonstrated that our method can generate variable-length motions, whose distribution is close to that of real motions.
4.1.2. Ablation studies
@@
Moreover, we discuss the effect of the dimension of latent space, adversarial training, and motion representation on the stage 1 model, and we utilize the reconstruction FID (rFID), MPJPE, FID, and MMDist as evaluation metrics. The dimensionality of the latent space may affect not only reconstruction quality but also influence the difficulty of training in stage 2, ultimately impacting the quality of the generated outputs. To investigate this, we evaluate the performance for cases with $d_{z}=\{8,16,32\}$. Table 2 shows that although the case of $d_{z}=16$ does not achieve the best generation quality compared to the other cases, it performs best in terms of FID and MMDist. Therefore, we adopted $d_{z}=16$ for MoLA. The results for VAE combined with GAN/SAN and modifying the encoder inputs are also shown in Table 2. As shown, the adversarial training is effective in the motion reconstruction task. It not only improves the reconstruction performance in stage 1 but also contributes to enhanced generation performance in stage 2. In particular, we can improve the performance of the stage 1 model by adopting the SAN framework instead of the conventional GAN. A better rFID is directly related to the upper bound of the overall performance of a text-to-motion model. Besides, the modifications to the encoder inputs in Equation (2) are shown to also have a comparable positive effect. Thus, we used VAE trained with the SAN framework (Section 3.2, Appendix) and the motion representation Equation (2) for our MoLA.
4.2. Motion Editing
Here, we demonstrate three different types of editing tasks with a single framework: editing with the path specified as the trajectory (path following), editing in the time direction (in-betweening), and editing in some joints of motion (upper body editing). In particular, for the path-following task, we quantitatively compare our approach with existing models (Xie et al., 2024; Dai et al., 2024).
Path-following: Path-following motion generation is a task of giving a trajectory (often the position of the pelvis) and generating the motion that matches the given route. Controlling the trajectory of generated motion enables more motion variation, avoidance of obstacles, and the creation of motion that meets physical constraints. Our framework can edit a motion to follow a specific path by minimizing the loss between a joint of the generated motion and a desired trajectory of the joint, applying the update rule of MPGD (He et al., 2024) on latent space (Section 3.4). In this experimental case, the desired pelvis trajectory is set as the control signal $\bm{y}$ in Equation (13). Figure 5 shows the editing results with our model when different path controls are given as $\bm{y}$ in the same text condition.
In addition, we show a quantitative comparison with existing methods: OmniControl (Xie et al., 2024) and MotionLCM (Dai et al., 2024), where the former is Continuous (raw data) type method and the latter is Continuous (raw latent) one. We compare them in terms of FID, R-Precision, Diversity, Trajectory err (50cm), Location err (50cm), Avgerage err, and Average inference time per sentence (AIST) in Table 3, following (Dai et al., 2024). Table 3 demonstrates that MoLA achieves competitive editing performance across all compared to other editing methods. Notably, while MoLA shows lower performance in following control signals compared to OminiControl, it achieves significantly faster editing. Note that OmniControl and MotionLCM have been training for this specific editing task, while our MoLA performs path-following without model fine-tuning. In other words, MoLA can perform other tasks without additional training as shown in the following sections, but OmniControl and MotionLCM require training a separate model for each task. MoLA is a more flexible and efficient framework.
In-betweening: Motion in-betweening is an important editing task that interpolates or fills the gaps between keyframes or major motion joints to create smooth 3D motion animation. Our framework can coordinate and generate motion between past and future contexts without additional training. We only need to set the start-end positions or the motions of a few frames as the control signal $\bm{y}$ in Equation (13). Figure 6 depicts the motion in-betweening results with our model when different start-end controls are given as $\bm{y}$ in the same text condition.
Upper body editing: Upper body editing combines generated upper body parts with given lower body parts. Generating some joints while keeping the other body joints following a given control signal can be seen as the task of outpainting in the spatial dimension of motion. In our framework, the loss and its gradient are calculated in the data space as Equation (12), and sampling is performed taking into account its gradients. The control signal $\bm{y}$ in Equation (13) is set to lower body positions that are not subject to editing. Figure 7 shows the upper body editing results with our model when different lower body controls are given as $\bm{y}$ in the same text condition. Note that, although we are dealing with upper body editing in this experiment, it is in principle possible to specify a different joint subset. Therefore, without the need for a separate architecture or additional training, our framework allows editing conditioned on given body parts of motion.
5. Conclusion
@@
In this paper, we proposed MoLA, a text-to-motion model that achieves fast, high-quality generation with multiple control tasks in a single framework. To construct our model, we rethought the representation of motion features and introduced an activation variable that characterizes the length of the motion. We also employed a latent diffusion-based method and introduced adversarial training. In addition, we applied a training-free guided generation framework based on MPGD. Our experiments showed that MoLA generates variable-length motions, whose distribution is close to that of real motions. Likewise, our quantitative evaluation demonstrated the effectiveness of adversarial training in text-to-motion generation. We also showed that our model can perform motion path-following, in-betweening, and upper body editing in a training-free manner.
@@ Appendix A SAN-based discriminator
Takida et al. (Takida et al., 2024) incorporated a sliced optimal transport perspective into a general GAN and established a new framework, slicing adversarial network (SAN). Following this work, we first decompose the discriminator $f_{\bm{\phi}}$ into the last linear layer and the remaining neural part, denoted as $\bm{w}\in\mathbb{R}^{d_{w}}$ and $h_{\bm{\varphi}}:\mathbb{R}^{N\times L}\to\mathbb{R}^{d_{w}}$ with $d_{w}\in\mathbb{N}$, respectively. This decomposition provides an interpretation of the discriminator: the neural function $h_{\bm{\varphi}}$ maps motion sequences to non-linear features, and then the linear layer $\bm{w}$ projects them into scalars. As a preliminary step for applying SAN to the adversarial training in Section 3.2.1, we normalize $w$ using its norm, resulting in $\bm{\omega}=\bm{w}/\|\bm{w}\|_{2}$, which indicates the direction of the projecting. Now, the discriminator is represented in the form of an inner product as $f_{\bm{\phi}}(\bm{x})=\bm{\omega}^{\top}h_{\varphi}(\bm{x})$, and its parameter is $\bm{\phi}=\{\bm{\varphi},\bm{\omega}\}$.
The prior work has shown that the discriminator obtained from the optimal solution of $\bm{\omega}$ in the hinge loss (5) does not guarantee gradients that make the generated distribution close to the data distribution. To address this issue, we adopt the SAN maximization problem instead of Equation (5). Specifically, we optimize the neural part $\bm{\varphi}$ with the original hinge loss, while applying the Wasserstein GAN loss (Arjovsky et al., 2017) to the direction $\bm{\omega}$. The modified maximization objective is formulated as
(14) $\mathcal{L}_{\text{SAN}}(\bm{\phi};\bm{\psi},\bm{\eta},\bm{x})$ $=\mathcal{L}_{\text{GAN}}(\{\bm{\varphi},\bm{\omega}^{-}\};\bm{\psi},\bm{\eta},\bm{x})+\bm{\omega}^{\top}\left(h_{\bm{\varphi}^{-}}(\bm{x})-\mathbb{E}_{q_{\bm{\eta}}(\bm{z}|\bm{x})}[h_{\bm{\varphi}^{-}}(g_{\bm{\psi}}(\bm{z}))]\right),$
@@ Appendix B Editing procedure on guided generation
In Algorithm 1, we show the specific procedure for guided generation described in Section 3.4. Note that we apply a time-travel technique to the update rule in Section 3.4 as in (Lugmayr et al., 2022; Wang et al., 2022; Yu et al., 2023; He et al., 2024) to achieve better editing results. This technique adds noise after each gradient descent step to implicitly perform a multi-step optimization of minimizing the distance measuring function $\mathfrak{L}(\cdot,\bm{y})$ and lead to an improvement in the editing quality.
@@ Appendix C Implementation Details
Training setup for stage 1 model: During the training, motion sequences are segmented into lengths of $L=64$. We use AdamW optimizer and batch size of 128. The hyperparameters in Eq. (4) and (6) are set as $\lambda_{\text{act}}=1.0$, $\lambda_{\text{reg}}=1.0\times 10^{-4}$, and $\lambda_{\text{adv}}=1.0\times 10^{-3}$. Our model $\{q_{\bm{\eta}},g_{\bm{\psi}},f_{\bm{\psi}}\}$ are trained with a simple multi-step learning late, the first 10,000 iterations with a learning rate of $2.0\times 10^{-4}$, and latter 5,000 with a learning rate of $2.0\times 10^{-5}$ in the case of HumanML3D dataset (Guo et al., 2022). Note that in Section E, we also conducted training and evaluation on a different dataset, the KIT-ML dataset (Plappert et al., 2016). For this dataset, the model was trained for the first 10,000 iterations with a learning rate of $5.0\times 10^{-5}$, followed by an additional 5,000 iterations with a reduced learning rate of $5.0\times 10^{-6}$. In addition, we replace the reconstruction loss in (4) with smooth $\ell_{1}$ loss (known as the special case of Huber loss) function and introduce a position enhancement term, following the technique in (Zhang et al., 2023b).

### v4

@@ 4 Experiments
We evaluate the performance of MoLA on two tasks: motion generation (Section 4.1) and motion editing (Section 4.2). Our results demonstrate that MoLA achieves its three key objectives: (1) fast and high-quality generation, (2) variable-length generation, and (3) multiple motion editing tasks in a training-free manner. To validate these objectives, we utilize the HumanML3D [12].333 This dataset contains 14,616 human motions from the AMASS [26] and HumanAct12 [11] datasets and 44,970 text descriptions.
4.1 Motion Generation
4.1.1 Performance comparison with other text-to-motion models
@@
We evaluate our proposed MoLA in comparison to current SOTA methods [21, 46, 37, 38, 48, 13, 24, 29, 44, 3, 49, 28, 6] using five metrics (R-Precision, Fréchet Inception Distance (FID), Multi-modal distance (MMDist), Diversity, and MultiModality (MModality)) proposed by Guo et al. [12]. For evaluation, we select the model that achieves the best FID, which is a metric that evaluates the overall motion quality, on the validation set and report its performance on the test set of HumanML3D. We show the results in Table 1. The methods are organized into three groups: i) those using VQ-based latent representations (Discrete), ii) those using data-space diffusion model (Continuous (raw data)), and iii) those using VAE-based latent representations (Continuous (latent)). The discrete approaches perform well in motion generation (e.g., [13, 24, 29]). However, those models cannot control an arbitrary set of joints in a training-free manner [29] as we have discussed so far. MDM [37], MotionDiffuse [46] and Fg-T2M [38] adopt data-space diffusion, while MLD [3], MotionLCM [6] and MoLA utilize a diffusion model in a lower-dimensional latent space. Therefore, the former are grouped in Continuous (raw data) and the latter in Continuous (latent) in the table. MoLA achieves the best performance in continuous methods, especially in the R-precision, FID, and MMDist metrics as shown in Table 1. Moreover, we also evaluate generation quality and inference cost in comparison to existing text-to-motion methods that have publicly available implementations [37, 3, 44, 29, 13]. Figure 2 shows that MoLA can generates much faster than MDM, which is the only existing method that performs training-free motion editing.
[…]
Editing type Methods
R-Precision
Top-3 $\uparrow$
Training-based editing OmniControl 0.688 0.192 9.533 0.065 0.007 0.053 74.4
MotionLCM 0.759 0.501 9.293 0.237 0.054 0.164 0.02
Table 2: Comparison of motion editing (path-following task) on HumanML3D dataset
4.1.2 Variable-length motion generation
Next, we discuss the impact of the activation variable in the motion representation introduced in Section 3.1. As shown in Equation (2), incorporating the activation variable into the motion features enables the Stage 2 model (Section 3.3) to generate variable-length motions. Figure 4 shows a comparison of the length distributions of samples generated by two existing methods (T2M-GPT and MoMask) and MoLA against the test set of HumanML3D. T2M-GPT represents a typical auto-regressive Text-to-Motion approach, while MoMask is a mask-prediction-based method that requires specifying the motion length. To address this limitation, MoMask introduces a length estimator that generates motion lengths conditioned on the input text. From Figure 4, we observe that our method successfully generates variable-length motions. Furthermore, the distribution of motion lengths generated by MoLA is closer to the actual motion distribution than those of the two existing methods, demonstrating the effectiveness of our approach.
4.2 Motion Editing
Here, we demonstrate three types of editing tasks using a unified framework: path-following (motion guided by a specified trajectory), in-betweening (editing in the time direction), and upper-body editing (modifying specific joints). In particular, for path-following, we quantitatively compare our approach with existing models [41, 6].
Figure 5: Qualitative results for the three editing tasks (path following, in-betweening, and upper-body editing). For these tasks, we treat each control signal (i) and (ii) in the left side of the figure as $\bm{y}$ in Equation (10) and (11). The corresponding generated results using the same input text are shown on the right side of the figure as (i) and (ii), respectively.
Path-following is a task of giving a trajectory (often the position of the pelvis) and generating the motion that matches the given route. Controlling the trajectory of generated motion enables more motion variation, avoidance of obstacles, and the creation of motion that meets physical constraints. In this experimental case, the desired pelvis trajectory is set as the control signal $\bm{y}$ in Equation (11). The upper row of Figure 5 shows the editing results with our model when different path controls are given as $\bm{y}$ in the same text condition. In addition, we show a quantitative comparison with existing methods: OmniControl [41] and MotionLCM [6], where the former is Continuous (raw data) type method and the latter is Continuous (raw latent) one. We compare them in terms of FID, R-Precision, Diversity, Trajectory err (50cm), Location err (50cm), Avgerage err, and Average inference time per sentence (AIST) in Table 2, following [6]. Table 2 demonstrates that MoLA achieves competitive editing performance across all compared to other editing methods. Notably, while MoLA shows lower performance in following control signals compared to OminiControl, it achieves significantly faster editing. Note that OmniControl and MotionLCM have been trained for this specific editing task, while MoLA performs path-following without model fine-tuning. In other words, MoLA can perform different tasks without additional training as shown in the following sections, but OmniControl and MotionLCM require training a separate model for each task. MoLA is a more flexible and efficient framework.
In-betweening is an important editing task that interpolates or fills the gaps between keyframes or major motion joints to create smooth 3D motion animation. Our framework can coordinate and generate motion between past and future contexts without additional training. We only need to set the start-end positions or the motions of a few frames as the control signal $\bm{y}$ in Equation (11). The middle row of Figure 5 depicts the motion in-betweening results with our model when different start-end controls are given as $\bm{y}$ in the same text condition.
Upper body editing combines generated upper body parts with given lower body parts. Generating some joints while keeping the other body joints following a given control signal can be seen as the task of outpainting in the spatial dimension of motion. The control signal $\bm{y}$ in Equation (11) is set to lower body positions that are not subject to editing. [^4] Note that, although we are dealing with upper body editing in this experiment, it is in principle possible to specify a different joint subset. The lower row of Figure 5 shows the upper body editing results with our model when different lower body controls are given as $\bm{y}$ in the same text condition.
@@ 4.2 Motion Editing
Table 3: Analysis of motion reconstruction and generation performance on HumanML3D dataset
4.3 Ablation studies
@@
We discuss the impact of latent space dimensionality, adversarial training, and motion representation on the stage 1 model. The dimensionality of the latent space may affect not only reconstruction quality but also influence the difficulty of training in stage 2, ultimately impacting the quality of the generated outputs. To investigate this, we evaluate the performance for cases with $d_{z}=\{8,16,32\}$. Table 3 shows that although the case of $d_{z}=16$ does not achieve the best reconstruction quality compared to the other cases, it performs best in terms of FID and MMDist. Therefore, we adopted $d_{z}=16$ for MoLA. The results for VAE combined with GAN/SAN and modifying the encoder inputs are also shown in Table 3. As shown in Table 3, the adversarial training is effective in the motion reconstruction task. It not only improves the reconstruction performance in stage 1 but also contributes to enhanced generation performance in stage 2. In particular, we can improve the performance of the stage 1 model by adopting the SAN framework instead of the conventional GAN. A better rFID is directly related to the upper bound of the overall performance of a text-to-motion model. Furthermore, as shown in Tables 3 and 4, modifying the encoder input Equation (2) improves reconstruction quality in terms of MPJPE without significantly compromising generation quality. As a result, this modification has a beneficial effect on the performance of the motion editing task, which requires fitting motion sequences into given control signals. This improvement can be attributed to the quality of the decoder $g_{\bm{\psi}}$, which contributes to the performance of the editing task, as indicated by Equations (10) and (11). Thus, we adopted VAE trained with the SAN framework [36] and the motion representation Equation (2) for MoLA.
Method Traj. err. $\downarrow$ Loc. err.$\downarrow$ Avg. err.$\downarrow$
Input for the encoder $q_{\bm{\eta}}$
$[(\bm{m}^{i})^{\mathsf{T}},a^{i}]^{\mathsf{T}}$ instead of Eq. (2) 0.281 0.068 0.174
MoLA ($d_{z}=16$) 0.271 0.051 0.159
Table 4: Analysis of motion editing (path-following task) performance on HumanML3D dataset
5 Conclusion
@@
We proposed MoLA, a text-to-motion model that achieves fast, high-quality generation with multiple control tasks in a single framework. We rethought the motion representation and introduced an activation variable that characterizes the length of the motion. In addition, we integrated latent diffusion, adversarial training, and a guided generation framework. Our experiments demonstrated MoLA’s ability to generate variable-length motions with distributions close to real motions and perform diverse motion editing tasks, significantly extending the performance boundaries of methods categorized as enabling training-free editing.
@@ Appendix A SAN-based discriminator
Takida et al. [36] incorporated a sliced optimal transport perspective into a general GAN and established a new framework, slicing adversarial network (SAN). Following this work, we first decompose the discriminator $f_{\bm{\phi}}$ into the last linear layer and the remaining neural part, denoted as $\bm{w}\in\mathbb{R}^{d_{w}}$ and $h_{\bm{\varphi}}:\mathbb{R}^{N\times L}\to\mathbb{R}^{d_{w}}$ with $d_{w}\in\mathbb{N}$, respectively. This decomposition provides an interpretation of the discriminator: the neural function $h_{\bm{\varphi}}$ maps motion sequences to non-linear features, and then the linear layer $\bm{w}$ projects them into scalars. As a preliminary step for applying SAN to the adversarial training in Section 3.2.1, we normalize $w$ using its norm, resulting in $\bm{\omega}=\bm{w}/\|\bm{w}\|_{2}$, which indicates the direction of the projecting. Now, the discriminator is represented in the form of an inner product as $f_{\bm{\phi}}(\bm{x})=\bm{\omega}^{\top}h_{\varphi}(\bm{x})$, and its parameter is $\bm{\phi}=\{\bm{\varphi},\bm{\omega}\}$.
The prior work has shown that the discriminator obtained from the optimal solution of $\bm{\omega}$ in the hinge loss (5) does not guarantee gradients that make the generated distribution close to the data distribution. To address this issue, we adopt the SAN maximization problem instead of Equation (5). Specifically, we optimize the neural part $\bm{\varphi}$ with the original hinge loss, while applying the Wasserstein GAN loss [1] to the direction $\bm{\omega}$. The modified maximization objective is formulated as
$\mathcal{L}_{\text{SAN}}(\bm{\phi};\bm{\psi},\bm{\eta},\bm{x})$ $=\mathcal{L}_{\text{GAN}}(\{\bm{\varphi},\bm{\omega}^{-}\};\bm{\psi},\bm{\eta},\bm{x})+\bm{\omega}^{\top}\left(h_{\bm{\varphi}^{-}}(\bm{x})-\mathbb{E}_{q_{\bm{\eta}}(\bm{z}|\bm{x})}[h_{\bm{\varphi}^{-}}(g_{\bm{\psi}}(\bm{z}))]\right),$ (12)
@@ Appendix B Editing procedure on guided generation
In Algorithm 1, we show the specific procedure for guided generation described in Section 3.4. Note that we apply a time-travel technique to the update rule in Section 3.4 as in [25, 39, 43, 14] to achieve better editing results. This technique adds noise after each gradient descent step to implicitly perform a multi-step optimization of minimizing the distance measuring function $\mathfrak{L}(\cdot,\bm{y})$ and lead to an improvement in the editing quality.
@@ Appendix C Implementation Details
Training setup for stage 1 model: During the training, motion sequences are segmented into lengths of $L=64$. We use AdamW optimizer and batch size of 128. The hyperparameters in Eq. (4) and (6) are set as $\lambda_{\text{act}}=1.0$, $\lambda_{\text{reg}}=1.0\times 10^{-4}$, and $\lambda_{\text{adv}}=1.0\times 10^{-3}$. Our model $\{q_{\bm{\eta}},g_{\bm{\psi}},f_{\bm{\psi}}\}$ are trained with a simple multi-step learning late, the first 10,000 iterations with a learning rate of $2.0\times 10^{-4}$, and latter 5,000 with a learning rate of $2.0\times 10^{-5}$ in the case of HumanML3D dataset [12]. Note that in Section E, we also conducted training and evaluation on a different dataset, the KIT-ML dataset [30]. For this dataset, the model was trained for the first 10,000 iterations with a learning rate of $5.0\times 10^{-5}$, followed by an additional 5,000 iterations with a reduced learning rate of $5.0\times 10^{-6}$. In addition, we replace the reconstruction loss in (4) with smooth $\ell_{1}$ loss (known as the special case of Huber loss) function and introduce a position enhancement term, following the technique in [44].